OpenStudy (anonymous):

Find the upper and lower limits of the zero of the following function. x^5-3x^3+24 plz explain, especially how to do the synthetic divison??

OpenStudy (amistre64):

5 4 3 2 1 0 1 0 -3 0 0 24 with synthetic we should line up our exponents with their coefficients

OpenStudy (amistre64):

this gives us a platform to work off of

OpenStudy (amistre64):

1 0 -3 0 0 24 0 add this row ----------------- zero | multiply this row

OpenStudy (amistre64):

if we choose a proper amount for our "zero" there will be no remainder on the end

OpenStudy (amistre64):

any questions so far?

OpenStudy (amistre64):

since the first coeff is a 1; lets get the factors of the last term: 24 1,24 2,12 3,8 4,6 and try them out ... as out zeros

OpenStudy (amistre64):

1 0 -3 0 0 24 0 2 4 2 4 8 ----------------- 2 | 1 2 1 2 4 32, not a zero and since they are all positive values we know that the solution aint greater than a 2 at least

OpenStudy (amistre64):

lets try a 1 1 0 -3 0 0 24 0 1 1 -2 -2 -2 ----------------- 1 | 1 1 -2 -2 -2 22 ; aint a 1

OpenStudy (amistre64):

lets try a -2 1 0 -3 0 0 24 0 -2 4 -2 4 -8 ----------------- -2| 1 -2 1 -2 4 16 , not a -2 it might not even be an integer value

OpenStudy (anonymous):

so what does that mean? that it's an x or something?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=x^5-3x^3%2B24

OpenStudy (amistre64):

it means that it gets ugly

OpenStudy (amistre64):

the graph shows that its got a root someplace between -2 and -2.5

OpenStudy (anonymous):

lol...why can math never be easy

OpenStudy (amistre64):

:) the world is to complicated for math to model it with ease

OpenStudy (anonymous):

lol I just wish sometimes my math didn't ask me to be psychic xD

OpenStudy (amistre64):

im not quite sure what it means by upper and lower limit tho

OpenStudy (anonymous):

I think it means the bounds

OpenStudy (anonymous):

the zero is greater than some number and less than some other number at the same time

OpenStudy (anonymous):

i have to find those two numbers somehow

OpenStudy (amistre64):

lets try a -3 1 0 -3 0 0 24 0 -3 9 -18 54 -162 ------------------- -3| 1 -3 6 -18 54 -138; well, its a + result at "-2" and a a - result at "-3" so if anything id say its in the interval: (-3,-2)

OpenStudy (amistre64):

a rather simple way to do this is just to try and err it till you get in the right ball park of the area

OpenStudy (amistre64):

f(x) = x^5 -3x^3 +24 f(-2) = = (-2)^5 -3(-2)^3 +24 = 16 f(-2.5) = (-2.5)^5 -3(-2.5)^3 +24 = -26.78125 its between -2 and -2.5 try -2.3 f(-2.3) = (-2.3)^5 -3(-2.3)^3 +24 = -3.862... still on the left of zero; go right a little f(-2.2) = (-2.2)^5 -3(-2.2)^3 +24 = 4.407 and "zero" in on it till youre close enough

OpenStudy (amistre64):

usually you have a "error" allowance so that you can provide a range; with a low and high boundary. without knowing the error allowance, its hard to say what a good interval will be

OpenStudy (anonymous):

I love how my math book does this to me. T.T

OpenStudy (anonymous):

thanks for the help

OpenStudy (amistre64):

yep, good luck :)